1.Order: chloral hydrate 350 mg po prior to electroencephalogram

Weight: 6.7 kg, infant

Supply: 500 mg/ 5ml

Literature: 25-50 mg/kg/dose po not to exceed 100 mg/kg

How many ml has been ordered for you to give? _____Is this a safe dose? YES or NO

2. Order: ceftazidime (Fortaz) 2 g IVPB q8h in 50 ml D51/2 NS

Weight: child 8 years, 24kg

Supply: 2g/10ml

Literature: safe dose is 2-6 g /24 hours given q8-q12 IV, concentration for IV 50mg/ml over 15-

30 minutes

Is this a safe dose? YES or NO

3. Order: Morphine 1.5 mg IV stat

Weight: 78 lb., child

Supply: 2mg/ml

Literature: 0.05 mg/kg/dose

How many ml will you give? _____Is this a safe dose? YES or NO

4. Order: chloral hydrate 200 mg po prior to electroencephalogram

Weight: 6.7 kg, infant

Supply: 500 mg/ 5ml

Literature: 25-50 mg/kg/dose po not to exceed 100 mg/kg

How many ml has been ordered for you to give? _____Is this a safe dose? YES or NO
Answer #1

ANSWER.

1) Calculation of volume(ml)of drug required.

     Formula

         Volume(ml) = Desired dose /Dose in hand * Quantity.

    Here

         Desired dose (D) = 350 mg.

         Dose in hand (H) = 500 mg.

             Quantity (Q) = 5 ml

   As per above formula

        Volume(ml) of drug required = D/H * Q

                                                      = 350/500 * 5

                                                      = 0.7 * 5

                                                      = 3.5 ml

       Safe dose = 25 to 50 mg/kg/dose not exceed 100 mg/kg.

       Maximum safe dose for a 6.7 kg child = 100 * 6.7

                                                    = 670 mg

         Ordered dose = 350 mg.

         * So the ordered dose is safe for this patient (YES)

2) Calculation of volume(ml) of drug required.

        Formula

           Volume(ml) = Desired dose/Dose in hand * Quantity.

      Here

          Desired dose (D) = 2 g q8h.

          Dose in hand (H) = 2 g

             Quantity (Q) = 10 ml

    As per above formula

        Volume(ml) of drug required = D/H * Q

                                                      = 2/2 * 10

                                                      = 10 ml

      Safe dose = 2g to 6 g /24 hrs.

     Ordered dose = 2 g (q8hr or three times a day).

                            = 2 * 3

                           = 6 g / 24 hrs.

    * So ordered dose is safe for this patient (YES).

3) Calculation of volume(ml) of drug required.

         Formula

           Volume(ml) = Desired dose/Dose in hand * Quantity.

    Here

        Desired dose(D) = 1.5 mg

        Dose in hand (H) = 2 mg

          Quantity (Q) = 1 ml

As per above formula

      Volume(ml) of drug required = D/H * Q

                                                    = 1.5/2 * 1

                                                    = 0.75 ml

        Safe dose = 0.05 mg/kg/dose.

   Weight of the patient = 78 lb.

     Convert lb to kg for calculation

           1 lb = 0.454 kg

   So

        Weight of the patient = 78 * 0.454

                                        = 35.4 kg.

    Safe dose for child weight 35.4 kg = 0.05 * 35.4

                                                            = 1.77 mg.

Ordered dose = 1.5 mg

* So ordered dose is safe for this patient (YES)

4) Calculation of volume(ml) of drug required.

        Formula

            Volume(ml) = Desired dose /Dose in hand * Quantity.

    Here

         Desired dose (D) = 200 mg.

         Dose in hand (H) = 500 mg

            Quantity (Q) = 5 ml

As per above formula

      Volume(ml) of drug required = D/H * Q

                                                    = 200/500 * 5

                                                    = 0.4 * 5

                                                    = 2 ml

     Safe dose = 25 to 50 mg/kg/dose and not to exceed 100 mg/kg.

     Maximum safe dose for a 6.7 kg child = 100 * 6.7

                                                  = 670 mg.

      Ordered dose = 200 mg

   * So ordered dose is safe for this patient (YES)